# The Feasibility of Topic-Based Watermarking on Academic Peer Reviews

### v1

@@ A.2 Fine-tuning Details
For instruction-tuned generation, we fine-tune the Llama-3.1-8B model using a parameter-efficient LoRA (Low-Rank Adaptation) method. LoRA freezes the original model weights and injects trainable low-rank matrices into a subset of layers, enabling effective fine-tuning with a small number of additional parameters. This approach is well-suited for large-scale models, reducing memory usage and training time while maintaining performance. Key settings include:
  • Adapter type: LoRA
  • LoRA $r$/$\alpha$: 16/32
  • LoRA Dropout: 0.1
  • Training epochs: 3
  • Batch size (per device): 2
  • Max sequence length: 2048 tokens
  • Learning rate: 1e-4
  • Warmup ratio: 0.2
  • Quantization: 4-bit (NF4), double quantization enabled
  • Target modules: q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj (These target modules correspond to the attention and MLP projections in transformer layers, where LoRA adapters are most effective.)
@@
All experiments were run using the Hugging Face Transformers and PEFT libraries, with training orchestrated using the Trainer API. The final adapters and tokenizer were saved for downstream evaluation. The dataset consists of the prompt (title, abstract, and generation instruction) and a completion (review text), compatible with instruction tuning for causal language models.
Appendix B Generation Quality Evaluations
@@
We expend our evaluation of topic-based watermarking (TBW) to assess its sensitivity to different token-to-topic similarity thresholds. In particular, we re-run perplexity and BERTScore evaluations using a lower semantic similarity threshold of $\tau=0.3$ (vs. $\tau=0.7$ in the main experiments). We also compare TBW against two baseline watermarking schemes, KGW and SynthID, to contextualize performance. We utilize and open-source watermarking framework, MarkLLM Pan et al. (2024), and the specified configurations for the baseline watermarking implementations.
B.1 Evaluation with Lower Topic Similarity Threshold ($\tau=0.3$)
@@
We repeat the perplexity and BERTScore evaluations described in Section 4.1.1 and Section 4.1.2 using a relaxed topic assignment threshold of $\tau=0.3$. This setting allows more tokens to be included in each green list, resulting in stronger watermark signals but potentially greater degradation in generation quality. The results help assess how sensitive TBW is to this design parameter.
B.1.1 Perplexity
@@
B.1.2 BERTScore Evaluation
We repeat the BERTScore F1 evaluation under the same setup described in Section 4.1.2, using generations produced with TBW at $\tau=0.3$. Results are shown in Figure 4.
@@ B.1.2 BERTScore Evaluation
B.2 Baseline Watermarking Quality
We compare TBW against two existing watermarking methods:
  • KGW Kirchenbauer et al. (2023): one of, if not the first watermarking approach for LLMs.
  • SynthID-Text (SynthID) Dathathri et al. (2024): Google’s proprietary watermarking technique designed for text attribution.
We evaluate their impact on fluency using perplexity and semantic similarity using BERTScore.
B.2.1 Perplexity
We evaluate perplexity for generations produced using KGW and SynthID, comparing their impact on fluency using the same evaluation framework as in Section 4.1.1. Figure 5 shows the perplexity distributions for each baseline, while Table 6 reports the number of samples with perplexity $\leq 20$ after truncation.
Figure 5: Perplexity distributions across model configurations with KGW and SynthID. Lower values indicate better fluency. Values above 20 are truncated for clarity.
Model Scheme Samples Retained
Base KGW 840
SynthID 538
Few-shot KGW 1000
SynthID 1000
Fine-tuned KGW 1000
SynthID 1000
Table 6: Number of retained generations with perplexity $\leq 20$ across model configurations, comparing KGW and SynthID.
Across all models, KGW performs reasonably well in preserving fluency. In the base model, its perplexity distribution is narrower and more favorable than that of SynthID, with 840 out of 1000 samples retained. In the few-shot setting, KGW is comparable to TBW at $\tau=0.7$, exhibiting slightly less variability. In the fine-tuned model, KGW performs better than TBW at $\tau=0.7$ and is similar in trend to TBW at $\tau=0.3$, suggesting its soft constraints are better tolerated by a model already adapted to the domain. In contrast, SynthID yields noticeably higher perplexity and wider distributions in the base and few-shot models, indicating reduced fluency and more frequent sampling of low-probability tokens. Only 538 base model generations were retained under the perplexity cap of 20. In the fine-tuned model, SynthID performs better, but still shows greater perplexity spread than KGW or TBW.
B.2.2 BERTScore Evaluation
We evaluate BERTScore F1 for generations produced with KGW and SynthID, using the same test setup and reference alignments as described in Section 4.1.2. Results are presented in Figure 6.
In the few-shot and fine-tuned configurations, KGW performs comparably to TBW at $\tau=0.7$, with similar median BERTScore values and distributional tightness. However, in the base model configuration, KGW shows a broader distribution of scores, indicating higher variability in semantic fidelity. This suggests that KGW, like TBW, is more effective when the generation is guided by conditioning or domain adaptation. SynthID shows a similar pattern but with slightly more pronounced effects. In the base model, SynthID outputs exhibit a wider spread compared to both TBW and KGW, reflecting less stable semantic alignment. In contrast, SynthID performs slightly better in the few-shot and fine-tuned settings, with a 1–2% improvement in BERTScore F1 over TBW at $\tau=0.7$.
Appendix C Robustness Evaluations
We provide additional details for the robustness evaluations described in Section 4.2. We include ROC curves for topic-based watermarking (TBW) and compare detection accuracy against the KGW and SynthID baselines under paraphrasing attacks. These results offer a more comprehensive view of how watermarking methods perform under realistic adversarial transformations.
C.1 ROC Curves
@@
Figure 7 presents ROC curves for TBW evaluated on outputs from the base, few-shot, and fine-tuned models. Detection performance remains strong in the base and fine-tuned settings, with area under the curve (AUC) values exceeding 0.90 under no attack and only moderately degraded under paraphrasing. The few-shot model is more sensitive to topic dilution, as discussed in Section 4.2, resulting in lower recall and reduced detection confidence under attack conditions.
C.2 Baseline Watermarking Robustness
To assess detection robustness of baseline methods, we apply the same paraphrasing attacks (PEGASUS and DIPPER) to generations produced by KGW and SynthID, and then evaluate each method’s ability to recover the watermark. Each row in Table 2 reflects detection accuracy out of 1,000 watermarked samples per setting.
Detection Accurcy
Language Model Attacks TBW KGW SynthID
Base No Attack 0.9460 0.9710 0.9090
PEGASUS 0.8470 0.4770 0.1350
DIPPER 0.8760 0.7540 0.1730
Few-shot No Attack 0.6220 0.9750 0.9590
PEGASUS 0.5800 0.5800 0.3590
DIPPER 0.5170 0.7480 0.2250
Fine-tuned No Attack 0.8800 0.9260 0.9600
PEGASUS 0.5830 0.4370 0.1800
DIPPER 0.5840 0.6570 0.1590
Table 7: Detection accuracy of TBW, KGW, and SynthID across model configurations and paraphrasing attack types. Each score reflects the proportion of correctly identified watermarked samples out of 1,000 examples per condition. Bolded values indicate the best result per row.
Under no-attack conditions KGW and SynthID outperform TBW in the few-shot and fine-tuned models. In the base model variant, TBW performs better than SynthID, but still worse than KGW with a smaller margin.
Under paraphrasing, TBW shows better robustness. In the base model, TBW outperforms KGW and SynthID by a wide margin, maintaining detection accuracy above 84% under PEGASUS and 87% under DIPPER. KGW degrades more sharply, and SynthID performs poorly across all paraphrasing conditions. In the few-shot setting, TBW and KGW perform similarly under PEGASUS, but TBW trails slightly under DIPPER. SynthID again suffers larger drops in accuracy. In the fine-tuned model, TBW maintains accuracy comparable to KGW and outperforming SynthID.
Appendix D Classifier Specifics
@@
We provide implementation details for the classification experiments described in Section 4.3.1. We outline the training setup used for both BERT and RoBERTa classifiers and summarize the evaluation strategy for attribution analysis on generated peer reviews.
D.1 Classifier Training
@@
Each model is fine-tuned using the Hugging Face Trainer API with early stopping based on F1. Key training settings include:
  • Model types: bert-base-uncased, roberta-large
  • Number of classes: 3 (reject, borderline, accept)
  • Max sequence length: 512 tokens
  • Training epochs: 5
  • Batch size (per device): 16
  • Learning rate: 2e-5
  • Warmup ratio: 0.1
  • Optimizer: AdamW
  • Scheduler: Cosine with restarts
  • Dropout: 0.2 (attention and hidden layers)
  • Gradient clipping: Max norm 1.0
  • Label smoothing: 0.1
  • Precision: Mixed (FP16 with full-eval)
  • Quantization: 4-bit weight loading (for memory efficiency)
  • Evaluation strategy: Per epoch; best model selected via F1 on validation set
  • Early stopping: Enabled (patience = 1)
@@ D.2 Classifier Evaluation
Predictions are obtained from each trained classifier on the tokenized test set and compared against the ground truth labels. We compute confusion matrices to visualize class-specific misclassification patterns and report overall accuracy as a coarse measure of performance. BERT achieves an accuracy of 51.3%, while RoBERTa performs slightly better at 53.9%. Figures 8 and 9 present the confusion matrices for BERT and RoBERTa, respectively.
@@
Figure 9: Confusion matrix for the RoBERTa classifier on on 1,000 human-written peer reviews.
@@
Appendix E Class-Specific Classifier Evaluation
To further characterize classifier performance, we conduct a class-specific evaluation of human-written peer reviews based on the same classification framework introduced in Section 4.3. This appendix extends the aggregate metrics reported in Table 3 by analyzing model behavior across the three target rating categories. Specifically, we examine confusion matrices for each classifier (BERT and RoBERTa), stratified by language model configuration (base, few-shot, fine-tuned) and watermarking condition (with or without topic-based watermarking). These matrices provide insight into the distribution of true versus predicted labels, allowing us to identify patterns of misclassification across rating levels.
Overall, we observe that classifier performance is strongest for the accept and borderline categories, with higher precision and recall scores relative to the reject class. This trend holds consistently across most configurations. The primary exception is observed in the BERT classifier applied to generations from the base LLM (without watermarking), where performance on the borderline class drops, leading to more frequent misclassifications into the neighboring categories.
This analysis underscores the relative semantic distinctiveness of strongly positive (Accept) and moderate (Borderline) reviews, while highlighting the challenges involved in distinguishing lower-quality (Reject) reviews, which often exhibit more linguistic and structural variability.
Appendix F Classifier-Based Attribution under Lower Topic Similarity Threshold ($\tau=0.3$)
We extend our classifier-based attribution analysis to topic-based watermarking (TBW) applied at a lower semantic similarity threshold of $\tau=0.3$, using the same evaluation methodology described in Section 4.3. This threshold relaxes the token-to-topic alignment constraints, thereby increasing green-list coverage and watermark signal strength, while potentially impacting semantic coherence.
Across classifiers and model variants, we observe a more balanced distribution of predictions among the three rating categories: accept, borderline, and reject. This suggests that the broader topic alignment may reduce overfitting to specific semantic patterns. However, in the fine-tuned model configuration, misclassifications of reject reviews remain more pronounced, indicating continued difficulty in capturing the linguistic signals associated with negative evaluations, even under stronger watermarking. The results are illustrated in Figure 10. Table 8 reports the classification metrics for each classifier and LLM model variant under TBW with $\tau=0.3$. While overall performance remains comparable to the $\tau=0.7$ condition, we observe that the fine-tuned model achieves the highest accuracy across both BERT and RoBERTa classifiers, suggesting that domain adaptation remains a dominant factor in attribution effectiveness even under relaxed topic alignment.
(a) BERT Base
(b) RoBERTa Base
(c) BERT Few-shot
(d) RoBERTa Few-shot
(e) BERT Fine-tuned
(f) RoBERTa Fine-tuned
Figure 10: Confusion matrices for topic-based watermarking (TBW) applied at a lower topic similarity threshold ($\tau=0.3$). Results are shown across all model configurations (base, few-shot, fine-tuned) and for both BERT and RoBERTa classifiers.

### v2

@@ C.3.2 Fine-Tuning Setup
For instruction-tuned generation, we fine-tune the Llama-3.1-8B model using a parameter-efficient LoRA (Low-Rank Adaptation) method. LoRA freezes the original model weights and injects trainable low-rank matrices into a subset of layers, enabling effective fine-tuning with a small number of additional parameters. This approach is well-suited for large-scale models, reducing memory usage and training time while maintaining performance. Key settings are described in Table 5.
Parameter Value
Adapter type LoRA
[…]
LoRA dropout 0.1
Training epochs 3
Batch size (per device) 2
Max sequence length 2048 tokens
Learning rate 1e-4
Warmup ratio 0.2
Quantization 4-bit (NF4), double quantization
Target modules q_proj, k_proj, v_proj,
o_proj, gate_proj, up_proj,
down_proj
Table 5: Fine-tuning Hyperparameters
@@
All experiments were conducted using the Hugging Face Transformers and PEFT libraries, with training orchestrated using the Trainer API. The final adapters and tokenizer were saved for downstream evaluation. The dataset consists of the prompt (title, abstract, and generation instruction) and a completion (review text), compatible with instruction tuning for causal language models.
Appendix D Generation Quality Evaluations
@@
We expand our evaluation of topic-based watermarking (TBW) to assess its sensitivity to different token-to-topic similarity thresholds. Specifically, we re-run perplexity and BERTScore evaluations using a lower semantic similarity threshold of $\tau=0.3$ (vs. $\tau=0.7$ in the main experiments). We also provide BERTScore comparisons of TBW ($\tau=0.7$) against KGW and SynthID baseline watermarking schemes.
D.1 Baseline Watermarking Parameters
We compare TBW against two established watermarking methods: KGW kirchenbauer2023watermark, a pioneering approach for LLM watermarking, and SynthID-Text dathathri2024scalable, Google’s proprietary technique for text attribution. All baseline implementations use the open-source MarkLLM framework pan-etal-2024-markllm with the following configurations:
KGW Parameters:
"gamma": 0.5,
"delta": 2.0,
"prefix_length": 1,
"z_threshold": 4.0
SynthID Parameters:
"ngram_len": 5,
"watermark_mode": non-distortionary,
"threshold": 0.52,
"detector_type": mean
D.2 BERTScore Evaluation (TBW $\tau=0.7$)
We use BERTScore F1 to evaluate semantic similarity between generated reviews and ground-truth references. This metric compares contextual embeddings and is tolerant to paraphrasing, making it well-suited for open-ended review generation. Results across all model configurations are shown in Figure 2.
D.3 Human Evaluation
To complement our automated text quality metrics, we conduct a small-scale human evaluation to assess the perceptual quality of TBW-generated peer reviews. Human evaluation provides crucial insights into whether watermark artifacts are detectable to end users, which is essential for practical deployment.
We conduct a single-set evaluation with 6 human evaluators to assess individual text samples across multiple quality dimensions. All evaluators came from a computer science background and were familiar with the review process of conferences used in our few-shot and fine-tuned model training (i.e., ICLR and NeurIPS).
Evaluators were presented with 15 generated review samples under our standard configuration ($\tau=0.7$, Llama-3.1-8B). For each sample, evaluators were shown the paper title, abstract, and generated review text. Of the 15 samples, 5 were generated using the base model, 5 using few-shot, and 5 using fine-tuned configurations. Sample order was randomized to avoid ordering effects. Evaluators rated each review on a 5-point Likert scale (1 = very poor, 5 = excellent) across three dimensions:
  • Fluency: “How fluent and natural is the text?”
  • Coherence: “How logically consistent and easy to follow is this text?”
  • Usefulness: “How useful would this review be for authors and reviewers?”
Table 6 summarizes the human evaluation results across all quality dimensions. As hypothesized, the base model performed worst across all metrics, with few-shot showing intermediate performance and fine-tuned achieving the highest scores. The fine-tuned model achieved mean ratings of 3.87 for fluency, 3.80 for coherence, and 3.60 for usefulness, demonstrating that domain adaptation via fine-tuning produces reviews perceived as substantially higher quality by human evaluators.
Model Fluency Coherence Usefulness
Base 2.80 $\pm$ 1.16 2.47 $\pm$ 0.97 2.07 $\pm$ 1.05
Few-shot 3.70 $\pm$ 0.65 3.47 $\pm$ 0.73 2.87 $\pm$ 1.07
Fine-tuned 3.87 $\pm$ 0.78 3.80 $\pm$ 0.76 3.60 $\pm$ 0.97
Table 6: Human evaluation results for TBW-generated reviews across model configurations. Scores are on a 5-point Likert scale (mean $\pm$ std. dev., $n=6$ evaluators, 5 samples per configuration).
One notable result is the usefulness rating of few-shot samples. While few-shot generation shows substantial improvements in fluency (3.70) and coherence (3.47) compared to the base model, usefulness remains relatively lower at 2.87. This pattern aligns with our findings in Section 4.2, where few-shot models exhibited degraded watermark detection performance under paraphrasing attacks. We hypothesize that topic misalignment between prompt exemplars and target papers weakens both topical consistency (affecting detection robustness) and review relevance (affecting perceived usefulness). Despite this limitation, the few-shot approach still produces reviews that human evaluators rate as reasonably fluent and coherent, suggesting that the watermarking mechanism itself does not introduce perceptible artifacts that degrade text quality.
D.4 BERTScore Evaluation (KGW & SynthID)
We evaluate BERTScore F1 for generations produced with KGW and SynthID. Results are presented in Figure 3.
In the few-shot and fine-tuned configurations, KGW performs comparably to TBW at $\tau=0.7$, with similar median BERTScore values and distributional tightness. However, in the base model configuration, KGW shows a broader distribution of scores, indicating higher variability in semantic fidelity. This suggests that KGW, like TBW, is more effective when the generation is guided by conditioning or domain adaptation. SynthID shows a similar pattern but with slightly more pronounced effects. In the base model, SynthID outputs exhibit a wider spread compared to both TBW and KGW, reflecting less stable semantic alignment. In contrast, SynthID performs slightly better in the few-shot and fine-tuned settings, with a 1-2% improvement in BERTScore F1 over TBW at $\tau=0.7$.
D.5 Evaluation with Lower Topic Similarity Threshold ($\tau=0.3$)
@@ D.4 BERTScore Evaluation (KGW & SynthID)
We repeat the perplexity and BERTScore evaluations using a relaxed topic assignment threshold of $\tau=0.3$. This setting allows more tokens to be included in each green list, resulting in stronger watermark signals but potentially greater degradation in generation quality. The results help assess how sensitive TBW is to this design parameter.
D.5.1 Perplexity
@@
D.5.2 BERTScore Evaluation
We repeat the BERTScore F1 evaluation, using generations produced with TBW at $\tau=0.3$. Results are shown in Figure 5.
@@ D.5.2 BERTScore Evaluation
Appendix E Robustness Evaluations
We provide comprehensive robustness analysis for the evaluations described in §4.2, including detailed performance metrics and ROC curve analysis across all experimental conditions.
@@ Appendix E Robustness Evaluations
Figure 6 presents ROC curves for TBW detection performance across base, few-shot, and fine-tuned model configurations under three conditions: no attack baseline, PEGASUS paraphrasing, and DIPPER paraphrasing. The curves demonstrate TBW’s consistent robustness across varying attack severity and model adaptation settings. Detection performance remains robust in base and fine-tuned configurations, with area under the curve (AUC) values exceeding 0.90 under no attack conditions and experiencing only moderate degradation under paraphrasing attacks. The few-shot model exhibits greater sensitivity to paraphrasing-induced topic dilution, as discussed in §4.2, resulting in reduced detection confidence and lower overall performance across all attack conditions.
Model Attack AUC F1 TPR@1% TPR@10%
Base None 0.968 0.955 0.908 0.956
PEGASUS 0.936 0.893 0.746 0.861
DIPPER 0.922 0.857 0.669 0.826
Few-shot None 0.729 0.768 0.626 0.669
PEGASUS 0.722 0.758 0.609 0.655
DIPPER 0.765 0.754 0.565 0.659
Fine-tuned None 0.981 0.927 0.817 0.948
PEGASUS 0.944 0.858 0.593 0.826
DIPPER 0.906 0.861 0.348 0.598
Table 8 provides comprehensive detection performance across all experimental conditions, including ROC-AUC, F1 scores, and true positive rates (TPR) at relevant false positive rate (FPR) thresholds of 1% and 10%. The base model demonstrates robustness, maintaining an AUC above 0.92 even under the strongest paraphrasing attacks, with F1 scores of 0.893 (PEGASUS) and 0.857 (DIPPER). At the 1% FPR threshold, the base model achieves TPR values of 74.6% and 66.9% under PEGASUS and DIPPER attacks, respectively, indicating strong practical utility for high-precision detection scenarios.
The fine-tuned model shows similar resilience with AUC values of 0.944 (PEGASUS) and 0.906 (DIPPER), though it exhibits greater sensitivity at low FPR thresholds, particularly under DIPPER attacks where TPR@1% drops to 34.8%. This suggests that fine-tuning may increase vulnerability to certain paraphrasing patterns while maintaining overall detection capability.
In contrast, the few-shot configuration shows limited degradation under attacks, with AUC values remaining stable around 0.72-0.77 across all conditions. However, the consistently lower baseline performance (AUC = 0.729) indicates that topic misalignment limits detectability in this setting, making the relative robustness less operationally significant.
The maintained performance at low FPR thresholds across model configurations confirms that TBW’s vocabulary partitioning strategy effectively preserves detection capability while minimizing false alarms on human-written content, as evidenced by the consistent TPR@1% and TPR@10% metrics across experimental conditions.
Appendix F Classifier Specifics
@@
We provide implementation details for the classification experiments described in §4.3. We outline the training setup used for both BERT and RoBERTa classifiers and summarize the evaluation strategy for attribution analysis on generated peer reviews.
F.1 Data Construction & Protocol
We first construct a labeled dataset by extracting review texts from our generation pipeline and assigning a class label based on the associated ground truth rating (e.g., scores 1-4 mapped to reject, 5-6 to borderline, and 7-10 to accept). To ensure accurate mapping, we align generated reviews with their original metadata using paper titles as unique identifiers. The final dataset consists of generated reviews paired with class labels, drawn from the fine-tuned generation split described in §C.3.2.
The final dataset consists of generated reviews paired with class labels, drawn from our experimental pipeline. The dataset is stratified into training and held-out test splits, with 9,000 balanced training samples (3,000 per class) and 1,000 test samples for evaluation.
F.2 Classifier Training
@@
Each model is fine-tuned using the Hugging Face Trainer API with early stopping based on F1. Key training settings described in Table 9.
Parameter Value
Model types bert-base-uncased,
roberta-large
Number of classes 3 (reject, borderline, accept)
Max sequence length 512 tokens
Training epochs 5
Batch size (per device) 16
Learning rate 2e-5
Warmup ratio 0.1
Optimizer AdamW
Scheduler Cosine with restarts
Dropout 0.2 (attention and hidden layers)
Gradient clipping Max norm 1.0
Label smoothing 0.1
Precision Mixed (FP16 with full-eval)
Quantization 4-bit weight loading
Evaluation strategy Per epoch;
best model selected via F1
Early stopping Enabled (patience = 1)
Table 9: Classifier Training Hyperparameters
@@ F.3 Classifier Evaluation
Predictions are obtained from each trained classifier on the tokenized test set and compared against the ground truth labels. We compute confusion matrices to visualize class-specific misclassification patterns and report overall accuracy as a coarse measure of performance. BERT achieves an accuracy of 51.3%, while RoBERTa performs slightly better at 53.9%. Figures 7 and 8 present the confusion matrices for BERT and RoBERTa, respectively.
@@
Figure 8: Confusion matrix for the RoBERTa classifier on 1,000 human-written peer reviews.
@@
F.4 Class-Specific Classifier Evaluation
To further characterize classifier performance, we conduct a comprehensive class-specific evaluation of generated peer reviews based on the classification framework introduced in §4.3. This analysis extends the aggregate metrics reported in Table 3 by examining model behavior across the three target rating categories under different watermarking conditions and topic similarity thresholds.
We examine confusion matrices for each classifier (BERT and RoBERTa), stratified by language model configuration (base, few-shot, fine-tuned) and watermarking condition. Additionally, we extend our analysis to topic-based watermarking (TBW) applied at a lower semantic similarity threshold of $\tau=0.3$, which relaxes the token-to-topic alignment constraints, thereby increasing green-list coverage and watermark signal strength while potentially impacting semantic coherence. Figure 9 presents the complete set of confusion matrices across all configurations, including both $\tau=0.7$ and $\tau=0.3$ conditions.
BERT exhibits distinct patterns across model configurations and watermarking conditions. In the base non-watermarked condition (panel a), BERT shows a skew toward the reject column, while the few-shot variant (panel b) demonstrates higher predictions in the accept and borderline columns. The fine-tuned non-watermarked model (panel c) shows the highest concentration in the borderline column, though values remain below 0.50. Under watermarking conditions, BERT base models with both TBW $\tau=0.7$ and $\tau=0.3$ (panels d, g) exhibit a slight skew toward the reject column but with modest values barely exceeding 0.50. For few-shot and fine-tuned watermarked variants (panels e, f, h, i), predictions concentrate in the borderline column, with accept predictions consistently higher than reject but lower than borderline.
RoBERTa demonstrates more consistent patterns with clearer biases toward specific categories. Across all base configurations, non-watermarked, TBW $\tau=0.7$, and TBW $\tau=0.3$ (panels j, m, p), there is a pronounced bias toward the borderline column. The few-shot and fine-tuned variants generally show better-balanced distributions with higher concentrations in both accept and borderline columns. Notable exceptions include the RoBERTa few-shot TBW $\tau=0.3$ condition (panel q), which maintains high borderline predictions, and the fine-tuned TBW $\tau=0.7$ variant (panel o), which also shows elevated borderline concentrations.